# Fast training and sampling of Restricted Boltzmann Machines

### v2

@@ Abstract
Restricted Boltzmann Machines (RBMs) are effective tools for modeling complex systems and deriving insights from data. However, training these models with highly structured data presents significant challenges due to the slow mixing characteristics of Markov Chain Monte Carlo (MCMC) processes. In this study, we build upon recent theoretical advancements in RBM training, focusing on the gradual encoding of data patterns into singular vectors of the coupling matrix, to significantly reduce the computational cost of training (in very clustered datasets) and evaluating and sampling in RBMs in general. The learning process is analogous to thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner. Such continuous transitions are associated with the critical slowdown effect, which adversely affects the accuracy of gradient estimates, particularly during the initial stages of training with clustered data. To mitigate this issue, we propose a pre-training phase that encodes the principal components into a low-rank RBM through a convex optimization process. This approach facilitates efficient static Monte Carlo sampling and accurate computation of the partition function. Furthermore, we exploit the continuous and smooth nature of the parameter annealing trajectory to achieve reliable and computationally efficient log-likelihood estimations, enabling online assessment during the training process, and proposing a novel sampling strategy termed parallel trajectory tempering that outperforms previously optimized MCMC methods. Our results demonstrate that this innovative training strategy enables RBMs to effectively address highly structured datasets that conventional methods struggle with. Additionally, we provide evidence that our log-likelihood estimation is more accurate than traditional, more computationally intensive approaches in controlled scenarios. Moreover, the parallel trajectory tempering algorithm significantly accelerates MCMC processes compared to existing and conventional methods.
@@ 1 Introduction
Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (Hinton & Sejnowski, 1983) and Restricted Boltzmann Machine (RBM) (Smolensky, 1986; Ackley et al., 1985) being among the earliest examples. They offer a clear framework for capturing complex data interactions and, with simple energy functions, can reveal underlying patterns, making them especially valuable for scientific applications. Pairwise-interacting models (Nguyen et al., 2017) have been widely used for decades for inference applications in fields like neuroscience (Hertz et al., 2011) and computational biology (Cocco et al., 2018). More recent research has shown that RBMs can also be reinterpreted as many-body physical models (Decelle et al., 2024; Bulso & Roudi, 2021), analyzed to identify hierarchical relationships in the data (Decelle et al., 2023), or used to uncover biologically relevant constituent features (Tubiana et al., 2019; Fernandez-de Cossio-Diaz, 2025). In essence, RBMs present a great compromise between expressibility and interpretability: they are universal approximators capable of modeling complex datasets and simple enough to allow a direct extraction of insights from data. This interpretative task is particularly challenging when dealing with more powerful models such as generative convolutional networks EBMs (Xie et al., 2016; Song & Kingma, 2021) and other state-of-the-art models in general. At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables, and where data is not abundant, which is often the case in genomics/proteomics or neural recording datasets. For instance, RBMs have shown to be among the most reliable approaches to generate synthetic human genomes (Yelmen et al., 2021; 2023) or to model inmunological response (Bravi et al., 2021a; b; 2023) or neural activity (van der Plas et al., 2023; Quiroz Monnens et al., 2024).
@@
Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general. The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes. Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects (Nijkamp et al., 2019; 2020; Decelle et al., 2021). This dynamical behavior can be explained analytically using moment-matching arguments  (Nijkamp et al., 2019; Agoritsas et al., 2023). While exploiting these effects can yield fast and accurate generative models, even for highly structured data (Carbone et al., 2023) or high-quality images with RBMs (Liao et al., 2022), this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters as investigated in Agoritsas et al. (2023); Decelle et al. (2024). Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
Both the ability to train an RBM in equilibrium and to generate convincing new samples from its equilibrium measure strongly depend on the dataset in question. For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps. However, this approach is no longer feasible for highly structured datasets (Béreux et al., 2023). Natural datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data, neural-recordings or low-temperature configurations in physical systems. These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries often associated to different labels. We show an example of the PCA of the 5 datasets we will be studying in this work in Fig. 1. The first 4 datasets are very clustered, while CelebA is not, as it presents a rather compact PCA. A very similar (non-clustered) situation is obtained if the entire MNIST dataset is considered. Details about these datasets are given in the caption of figure 1. During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution. Sampling from multimodal distributions is particularly challenging because the mixing times are controlled by the transition times between modes. But this is not the only difficulty. In RBMs, these distant modes can be shown analytically(Decelle et al., 2017; 2018; Bachtis et al., 2024) to gradually appear in the model’s equilibrium measure during training by a process formally analogous to the ferromagnetic phase transition in simple spin models, the process by which the magnetic spins of a material spontaneously align with a fixed pattern when the temperature goes below a certain transition temperature. The RBM training is characterized as a sequence of such phase transitions in which the specialization of the model parameters plays the role of temperature reduction. These pattern-encoding phase transitions are continuous phase transitions (also known as second-order transitions or critical transitions), and these types of transitions are associated with a sudden arrest in the characteristic times of the visible variables, a phenomenon known as critical slowdown, which causes the MCMC mixing time to diverge with a power of the system size when the transition point is approached. The critical slowdown effects that occur in the training trajectory were recently studied in Bachtis et al. (2024). A brief introduction to phase transitions and the associated dynamical effects can be found in the SI G.
This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples. Obtaining new and independent configurations would require an impractically large number of sampling steps.
@@ 2 Related work
Training EBMs by maximizing log-likelihood has long been a challenge in the community (LeCun et al., 2006; Song & Kingma, 2021). EBMs gained popularity with the introduction of the contrastive divergence algorithm (Hinton, 2002), in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps. Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models (Salakhutdinov & Murray, 2008; Desjardins et al., 2010; Decelle et al., 2021). An improvement is the persistent contrastive divergence (PCD) algorithm (Tieleman, 2008), which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused. PCD acts like a slow annealing process improving gradient estimation quality. However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model (Béreux et al., 2023). This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified. For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions (Béreux et al., 2023). Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties. Other optimized MCMC methods, such as the Parallel Tempering (PT) algorithm  (Hukushima & Nemoto, 1996) (in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule), have been proposed to improve the training process (Salakhutdinov, 2009; Desjardins et al., 2010) and show important improvements in certain datasets (Krause et al., 2020).
@@
However, PT is expensive because it requires sampling at multiple temperatures while using samples from only one model. It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by Decelle & Furtlehner (2021a). This leads to sharp drops in acceptance rates for exchanges between temperatures near the transition. For this reason, first-order transitions typically hinder PT, as many intermediate temperatures are needed to allow sufficient acceptance rates in the exchanges. In contrast, continuous transitions, as seen during the training trajectory, cause modes to gradually separate as the model parameters are changed.
The population annealing algorithm, which adjusts the statistical weights of parallel chains during learning each time the parameters are updated, has been proposed as an alternative approach (Krause et al., 2018). Reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has also been explored (Carbone et al., 2024). However, both methods struggle with highly structured data, as frequent reweighting (which involves duplicating only a few samples) biases the equilibrium measure due to an insufficient number of effective independent chains. To prevent chains from becoming too correlated around transition regions during training, either more sampling steps or a lower learning rate is required, which leads to very long training times to ensure equilibrium. In our case, we tried to used both methods with our datasets but the training becomes too unstable or two slow if learning rate and number of Gibbs steps are auto adjusted. Another approach uses EBMs as corrections for flow-based models that are easier to sample (Nijkamp et al., 2022), simplifying learning but sacrificing the interpretability of the energy function, a key goal of this work. Alternatively, evolving flow models can serve as fast sampling proposers for the EBM (Grenioux et al., 2023), but this requires training two separate networks and the acceptance of the proposed moves is significantly reduced as the EBM becomes more specialized.
@@
A recent method for RBMs, called “stacked tempering” (Roussel et al., 2023), significantly accelerates sampling by training smaller RBMs using latent variables from previous models. This allows for fast updates through a PT-like algorithm, and the authors demonstrated that it is much faster than standard PT. However, while this method is effective for sampling a trained model, it is not suitable for learning, as it requires sequentially training multiple models stacked on top of each other. Also for RBMs, it has recently been shown in a theoretical work that it should be possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see Decelle & Furtlehner (2021a) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine. This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process. While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension, and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
@@
In this paper, we show how starting the RBM training with a low-rank RBM in very clustered datasets can drastically improve the quality of the models by circumventing the initial dynamical arrest effects that occur during the training process, at a very similar computational cost. We also show that the training trajectory can be leveraged to estimate the log-likelihood at a very little computational effort, so that both the training and test set log-likelihood can be easily and reliably computed online during training. Furthermore, we propose an efficient sampling method, which we call Parallel Trajectory Tempering (PTT). This method outperforms optimized MCMC methods such as the traditional Parallel Tempering (PT) or the “Stacked Tempering” (Roussel et al., 2023) only requiring storing the parameters of the model at a few stages during training. By combining these strategies, we are able to train equilibrium models that accurately capture different modes in datasets prone to mode collapse using standard methods, easily evaluate RBMs during the training process, and easily sample from multimodal distribution.
@@ 3 The Restricted Boltzmann Machine
The RBM is composed by $N_{\mathrm{v}}$ visible nodes and $N_{\mathrm{h}}$ hidden nodes. In our study, we primarily use binary variables $\{0,1\}$ or ±$\pm 1$ for both layers. The two layers (visible and hidden) interact via a weight matrix $\bm{W}$, with no direct couplings within a given layer. Variables are also adjusted by visible and hidden local biases, $\bm{\theta}$ and $\bm{\eta}$, respectively. The Gibbs-Boltzmann distribution for this model is expressed as
@@
$\frac{\partial\mathcal{L}}{\partial w_{ia}}$ $=\langle v_{i}h_{a}\rangle_{\mathcal{D}}-\langle v_{i}h_{a}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\langle v_{i}\rangle_{\mathcal{D}}-\langle v_{i}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\eta_{a}}=\langle h_{a}\rangle_{\mathcal{D}}-\langle h_{a}\rangle_{\rm RBM},$ (2)

### v3

@@ Abstract
Restricted Boltzmann Machines (RBMs) are powerful tools for modeling complex systems and extracting insights from data, but their training is hindered by the slow mixing of Markov Chain Monte Carlo (MCMC) processes, especially with highly structured datasets. In this study, we build on recent theoretical advances in RBM training and focus on the stepwise encoding of data patterns into singular vectors of the coupling matrix, significantly reducing the cost of generating new samples and evaluating the quality of the model, as well as the training cost in highly clustered datasets. The learning process is analogous to the thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner. We leverage the continuous transitions in the training process to define a smooth annealing trajectory that enables reliable and computationally efficient log-likelihood estimates. This approach enables online assessment during training and introduces a novel sampling strategy called Parallel Trajectory Tempering (PTT) that outperforms previously optimized MCMC methods. To mitigate the critical slowdown effect in the early stages of training, we propose a pre-training phase. In this phase, the principal components are encoded into a low-rank RBM through a convex optimization process, facilitating efficient static Monte Carlo sampling and accurate computation of the partition function. Our results demonstrate that this pre-training strategy allows RBMs to efficiently handle highly structured datasets where conventional methods fail. Additionally, our log-likelihood estimation outperforms computationally intensive approaches in controlled scenarios, while the PTT algorithm significantly accelerates MCMC processes compared to conventional methods.
@@ 1 Introduction
Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (hinton1983optimal) and Restricted Boltzmann Machine (RBM) (Smolensky; ackley1985learning) being among the earliest examples. They offer a clear framework for capturing complex data interactions and, with simple energy functions, can reveal underlying patterns, making them especially valuable for scientific applications. Pairwise-interacting models (nguyen2017inverse) have been widely used for decades for inference applications in fields like neuroscience (hertz2011ising) and computational biology (cocco2018inverse). More recent research has shown that RBMs can also be reinterpreted as many-body physical models (decelle2024inferring; bulso2021restricted), analyzed to identify hierarchical relationships in the data (decelle2023unsupervised), or used to uncover biologically relevant constituent features (tubiana2019learning; Fernandez-de-Cossio-Diaz2025). In essence, RBMs present a great compromise between expressibility and interpretability: they are universal approximators capable of modeling complex datasets and simple enough to allow a direct extraction of insights from data. This interpretative task is particularly challenging when dealing with more powerful models, such as generative convolutional network EBMs (pmlr-v48-xiec16; song2021train) and other state-of-the-art models in general. At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables and where data are not abundant, which is often the case in genomics/proteomics or neural recording datasets. For example, RBMs have been shown to be among the most reliable approaches to generate synthetic human genomes (yelmen2021creating; yelmen2023deep) or to model the immune response (bravi2021probing; bravi2021rbm; bravi2023transfer) or neural activity (van2023neural; Quiroz_Monnens_2024).
@@
RBMs, like other EBMs, are challenging to train due to the difficulty of computing the log-likelihood gradient, which requires ergodic exploration of a complex free-energy landscape via Markov Chain Monte Carlo (MCMC). Training with non-convergent MCMC introduces out-of-equilibrium memory effects, as shown in recent studies  (nijkamp2019learning; nijkamp2020anatomy; decelle2021equilibrium), which can be analytically explained using moment-matching arguments  (nijkamp2019learning; agoritsas2023explaining). While these effects enable fast and accurate generative models for structured data (carbone2024fast) and high-quality images with RBMs (liao2022gaussian), they create a sharp disconnect between the model’s Gibbs-Boltzmann and dataset distributions, undermining parameter interpretability (agoritsas2023explaining; decelle2024inferring). For meaningful insights, RBMs require proper chain mixing during training, emphasizing the need for equilibrium models.
Working with equilibrium RBMs poses three primary challenges: training, evaluating model quality, and sampling from the trained model. Both training and sampling rely on MCMC methods to draw equilibrium samples, but their efficiency depends heavily on the dataset’s structure. While RBMs can be effectively trained on image datasets like MNIST or CIFAR-10 with sufficient MCMC steps, this approach falters for highly structured datasets (bereux2023learning), such as genomics, neural recordings, or low-temperature physical systems. These datasets often exhibit multimodal distributions with distinct clusters, as revealed by PCA (Fig. 1), which hinder mixing. Clustering is prominent in the first four datasets studied here, in contrast to the more compact PCA pattern seen in CelebA or full MNIST. Dataset details are provided in Fig. 1. Multimodal distributions exacerbate sampling challenges since mixing times depend on the MCMC chains’ ability to transition between clusters. In RBMs, these distant modes emerge during training through continuous phase transitions (e.g., second-order or critical transitions) that encode patterns (Decelle_2017; decelle2018thermodynamics; bachtis2024cascade). These transitions lead to critical slowdown (hohenberg1977theory), where mixing times diverge with system size as the transition point nears. Critical slowdown along the training trajectory was recently analyzed in (bachtis2024cascade). For a brief introduction to phase transitions and associated dynamical effects, see Supplemental Information (SI) G. All this hampers both the training process and the model’s ability to generate independent samples, as achieving adequate mixing requires prohibitively long simulations. Model quality evaluation adds to the challenge, particularly in unsupervised learning. A common qualitative approach compares equilibrium samples to data via PCA projections. Quantitatively, moment comparisons between data and generated distributions offer insights but fail to detect mode collapse. Log-likelihood computation provides a single-score metric and tracks overfitting during training, but exact computation is infeasible due to the intractable partition function. Instead, methods such as Annealed Importance Sampling (AIS) (neal2001annealed) are used for estimation but become unreliable for highly multimodal distributions and are computationally impractical for online evaluations (see Section 4.1).
In this paper, we build on recent advances in understanding the evolution of the landscape during the training of RBM models to propose 3 innovative approaches for tackling the challenges of evaluating, generating samples and training RBMs on highly structured datasets:
  1. 1.
    We propose a novel method for estimating log-likelihood (LL) by leveraging the softness of the training trajectory. This approach enables efficient online LL computation during training at minimal cost or offline reconstruction using trajectory annealing or parallel tempering. Moreover, we demonstrate that this method is not only less computationally expensive than existing techniques but also more accurate in controlled experiments. The method is introduced in Sect. 4.1.
  2. 2.
    We propose a variation of the standard parallel tempering algorithm, where parameters from different stages of the learning process are exchanged instead of temperatures. This new sampling approach significantly outperforms previous optimized methods and is detailed in Section 4.2.
  3. 3.
    We introduce a pretraining strategy based on the low-rank RBM framework of (decelle2021exact) to mitigate gradient degradation in early training, caused by mixing time divergence when encoding distant modes in structured datasets. Our approach extends this framework to real datasets by increasing constrained directions and incorporating a trainable bias, essential for image data. We also refine static Monte Carlo sampling for efficient equilibrium sampling. See Section 5 and SI A for details.
We also provide an in-depth discussion in Sect. 2 on the physical reasons behind the failure of previous algorithms when handling clustered datasets.
@@ 2 Related work
Training EBMs by maximizing log-likelihood has long been a challenge (lecun2006tutorial; song2021train). The introduction of contrastive divergence (hinton2002training) popularized EBMs by using short MCMC runs initialized on minibatch examples, but this approach often produces models with poor equilibrium properties, limiting their effectiveness as generative models (salakhutdinov2008quantitative; desjardins2010tempered; decelle2021equilibrium). Persistent contrastive divergence (PCD) (tieleman2008training) improves gradient estimation by maintaining a continuous MCMC chain between updates, acting like a slow annealing process. However, PCD struggles with clustered data, as the chain’s properties can drift from the equilibrium measure, degrading the model (bereux2023learning). This issue, tied to phase coexistence, can be mitigated with constrained MCMC methods, leveraging order parameters from the singular value decomposition of the RBM coupling matrix to reconstruct multimodal distributions (bereux2023learning). While effective for model evaluation, this approach is too computationally intensive for practical training, despite producing models with strong equilibrium properties.
The population annealing algorithm, which adjusts the statistical weights of parallel chains after each parameter update, has been proposed as an alternative (krause2018population). Similarly, reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has been explored (carbone2024efficient). However, both methods face challenges with highly structured data, where frequent reweighting (duplicating only a few samples) skews the equilibrium measure due to an insufficient number of independent chains. To mitigate this, more sampling steps or lower learning rates are required, leading to excessively long training times to maintain equilibrium.
In our case, applying these methods to our datasets resulted in training that was either too unstable or prohibitively slow when auto-adjusting learning rates and Gibbs steps. Another approach uses EBMs as corrections for flow-based models (nijkamp2022mcmc), simplifying sampling but sacrificing the interpretability of the energy function, which is central to our work. Alternatively, evolving flow models can act as fast sampling proposers for EBMs (grenioux2023balanced), but this requires training two separate networks and leads to lower acceptance rates as the EBM becomes more specialized.
The Parallel Tempering (PT) algorithm  (hukushima1996exchange), in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule, has been proposed to both speed up the sampling and to improve the training process (salakhutdinov2009learning; desjardins2010tempered) showing important improvements in certain datasets (krause_algorithms_2020).
@@
However, PT is expensive because it requires sampling at multiple temperatures while using samples from only one model. It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by (decelle2021exact). This leads to sharp drops in acceptance rates for exchanges between temperatures near the transition. For this reason, first-order transitions typically hinder PT, as many intermediate temperatures are needed to allow sufficient acceptance rates in the exchanges. In contrast, continuous transitions, as seen during the training trajectory, cause modes to gradually separate as the model parameters are changed. The “Stacked Tempering” algorithm (roussel2023accelerated), recently proposed for RBMs, accelerates sampling by training smaller RBMs with latent variables from previous models, enabling fast updates via a PT-like algorithm. While significantly faster than PT, it is unsuitable for learning, as it requires sequential training of multiple stacked models.
@@
Separately, a theoretical study (decelle2021exact) demonstrated that a low-rank RBM can be trained via a convex, fast optimization process to reproduce the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see (decelle2021exact) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine. This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process. While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
@@
In this paper, we show that the training trajectory can be exploited to estimate log-likelihood with minimal computational cost, enabling reliable online computation. We also introduce Parallel Trajectory Tempering (PTT), an efficient sampling method that outperforms optimized MCMC approaches like Parallel Tempering (PT) and Stacked Tempering (roussel2023accelerated), requiring only a few model parameters stored at different stages of the training. Additionally, we demonstrate that initializing RBM training with a low-rank RBM in highly clustered datasets mitigates the effect of initial dynamical arrests, significantly improving model quality at a comparable cost. Combining these strategies allows us to train equilibrium models that capture diverse modes in datasets prone to mode collapse, efficiently evaluate RBMs during training, and sample from multimodal distributions.
@@ 3 The Restricted Boltzmann Machine
The RBM consists of $N_{\mathrm{v}}$ visible nodes and $N_{\mathrm{h}}$ hidden nodes. In our study, we primarily use binary variables $\{0,1\}$ or $\pm 1$ for both layers. The visible and hidden layers interact through a weight matrix $\bm{W}$, with no intra-layer couplings. Each variable is further influenced by local biases, $\bm{\theta}$ for visible units and $\bm{\eta}$ for hidden units. The Gibbs-Boltzmann distribution for this model is given by
@@
$\textstyle\frac{\partial\mathcal{L}}{\partial W_{ia}}$ $=\langle v_{i}h_{a}\rangle_{\mathcal{D}}-\langle v_{i}h_{a}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\langle v_{i}\rangle_{\mathcal{D}}-\langle v_{i}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\eta_{a}}=\langle h_{a}\rangle_{\mathcal{D}}-\langle h_{a}\rangle_{\rm RBM},$ (2)
where $\langle f(\bm{v},\bm{h})\rangle_{\mathcal{D}}=M^{-1}\sum_{m}\sum_{\{\bm{h}\}}f(\bm{v}^{(m)},\bm{h})p(\bm{h}|\bm{v}^{(m)})$ denotes the average with respect to the entries in the dataset, and $\left\langle{f(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$.